# A Mixture of Experts Vision Transformer for High-Fidelity Surface Code Decoding

### v1

@@ 3.2 Toric Code
A particularly important stabilizer code family is the toric code, which provides a prototypical example of a two dimensional topological stabilizer code. It is defined on a $L\times L$ square lattice with periodic boundary conditions, topologically a torus, with physical qubits placed on edges. We write $e$ for an edge of the lattice, $\delta(v)$ for the set of edges incident to a vertex $v$, and $\partial p$ for the set of boundary edges of a plaquette (face) $p$. The stabilizer generators are local Pauli operators: star checks at vertices,
@@ 3.3 Mixture of Experts
First introduced in Jacobs et al. (1991), the Mixture of Experts (MoE) framework is an ensemble architecture that combines multiple specialized submodels through an adaptive gating mechanism. In a typical MoE layer, let there be $n$ expert $f_{1},\ldots,f_{n}$, and a sequence of input tokens represented by $\mathbf{X}=[x_{1},\ldots,x_{N}]\in\mathbb{R}^{N\times d}$, where $N$ is the number of tokens in the sequence. Consequently, the output of the dense MoE layer can be reformulated as
@@ 4.2 QuantumSMoE
where the weighting coefficients $\lambda_{\text{BER}}$, $\lambda_{\text{LER}}$ are selected by $0.5$ and $1$ as the setting in Choukroun and Wolf (2024), while $\lambda_{\text{os}}$ is set to $0.1$ via grid search to achieve optimal performance.
@@ 5.1 Experimental Setting
Implementation Details. Our model is trained using 6 layers, with MoE layers applied in the final two blocks to reduce decoding latency. We set the embedding dimension to 128, following the same design adopted in QECCT. The SoftMoE layer is configured with 8 experts, where each expert processes 4 slots using a feedforward network with a hidden dimension of 512. The number of attention heads is set to $8$. For training, we use $5,760,000$ samples generated from nine physical error rates ranging from 0.05 to 0.2. The model is trained for 200 epochs using the AdamW optimizer, with an initial learning rate of 0.001 and a cosine decay schedule down to $10^{-6}$. Training is performed on a single NVIDIA RTX 4090 GPU with an AMD EPYC 7282 CPU (16 cores).
@@ 5.2 Main Results
Next, we examine the impact of ViT based architecture and MoE integration in our decoder. Figure 5 evaluates the efficacy of the MoE block within our quantum decoder. Since our model utilizes a ViT backbone, it inherently incorporates local information through patch embeddings. We further compare our approach with QECCT to demonstrate how the inclusion of patch embeddings provides a crucial spatial inductive bias. Unlike the standard Transformer framework of QECCT, which processes syndromes without explicit localized context, our model leverages these embeddings to effectively capture the underlying connectivity between qubits. Figure 5 illustrates that even without the MoE component, QuantumSMoE consistently outperforms QECCT for distances $L=6$ and $8$. These findings suggest that incorporating spatial information is crucial for the model to learn the underlying correlations between errors. Notably, the addition of MoE layers provides a significant performance boost with very little overhead on computational cost. This integration enables the model to achieve decoding performance that consistently outperforms prior learning based decoders under the considered settings.

### v2

@@ 3.2 Toric Code
A particularly important stabilizer code family is the toric code, which provides a prototypical example of a two dimensional topological stabilizer code. It is defined on a $L\times L$ square lattice with periodic boundary conditions, topologically a torus, with physical qubits placed on edges. We write $e$ for an edge of the lattice, $\delta(v)$ for the set of edges incident to a vertex $v$, and $\partial p$ for the set of boundary edges of a plaquette (face) pp. The stabilizer generators are local Pauli operators: star checks at vertices,
@@ 3.3 Mixture of Experts
First introduced in Jacobs et al. (1991), the Mixture of Experts (MoE) framework is an ensemble architecture that combines multiple specialized submodels through an adaptive gating mechanism. In a typical MoE layer, let there be $n$ expert $f_{1},\ldots,f_{n}$, and a sequence of input tokens represented by $\mathbf{X}=[x_{1},\ldots,x_{N}]\in\mathbb{R}^{N\times d}$, where NN is the number of tokens in the sequence. Consequently, the output of the dense MoE layer can be reformulated as
@@ 4.2 QuantumSMoE
where the weighting coefficients $\lambda_{\text{BER}}$, $\lambda_{\text{LER}}$ are selected by $0.5$ and 11 as the setting in Choukroun and Wolf (2024), while $\lambda_{\text{os}}$ is set to $0.1$ via grid search to achieve optimal performance.
@@ 5.1 Experimental Setting
Implementation Details. Our model is trained using 6 layers, with MoE layers applied in the final two blocks to reduce decoding latency. We set the embedding dimension to 128, following the same design adopted in QECCT. The SoftMoE layer is configured with 8 experts, where each expert processes 4 slots using a feedforward network with a hidden dimension of 512. The number of attention heads is set to 88. For training, we use $5,760,000$ samples generated from nine physical error rates ranging from 0.05 to 0.2. The model is trained for 200 epochs using the AdamW optimizer, with an initial learning rate of 0.001 and a cosine decay schedule down to $10^{-6}$. Training is performed on a single NVIDIA RTX 4090 GPU with an AMD EPYC 7282 CPU (16 cores).
@@ 5.2 Main Results
Next, we examine the impact of ViT based architecture and MoE integration in our decoder. Figure 5 evaluates the efficacy of the MoE block within our quantum decoder. Since our model utilizes a ViT backbone, it inherently incorporates local information through patch embeddings. We further compare our approach with QECCT to demonstrate how the inclusion of patch embeddings provides a crucial spatial inductive bias. Unlike the standard Transformer framework of QECCT, which processes syndromes without explicit localized context, our model leverages these embeddings to effectively capture the underlying connectivity between qubits. Figure 5 illustrates that even without the MoE component, QuantumSMoE consistently outperforms QECCT for distances $L=6$ and 88. These findings suggest that incorporating spatial information is crucial for the model to learn the underlying correlations between errors. Notably, the addition of MoE layers provides a significant performance boost with very little overhead on computational cost. This integration enables the model to achieve decoding performance that consistently outperforms prior learning based decoders under the considered settings.